# Parsimonious Compactly Supported Covariance Models in the Gauss Hypergeometric Class: Identifiability, Reparameterizations, and Asymptotic Properties

### v1

@@ 4 Fixing the identifiability problem: the Hypergeometric correlation model
In this Section we first provide necessary and sufficient conditions for the validity of the ${\cal GH}$ model. Then we solve the identifiability problem by considering a parsimonious version that responds to an optimality criterion. With this goal in mind, we consider a ${\cal GW}$-type reparameterization of the $\mathcal{GH}_{\delta,\beta,\gamma,a}$ model, that is, we set
$\delta=\frac{d+1}{2}+\kappa,\quad\beta=\delta+\frac{\mu}{2},\quad\gamma=\beta+l,$ (15)
with $l\geq 0$. As outlined in (13), this parameterization includes the ${\cal GW}$ model as a special case when $l=\frac{1}{2}$.
The following theorem can be viewed as an improvement of Theorem 1 in Emery and Alegría, (2022) and provides necessary and sufficient conditions for the validity of the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ model.
@@
Let $d$ be a fixed dimension, $\delta=\frac{d+1}{2}+\kappa$ and $\kappa>-\frac{1}{2}$ a smoothness parameter. Then $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}$ if and only if
@@
    $\mu\geq\delta-l+\frac{1}{2}\quad$ if $\quad 0\leq l\leq\frac{d}{2}+\kappa$
@@
    $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l\quad$ if $\quad l>\frac{d}{2}+\kappa$.
It should be stressed that necessary and sufficient conditions can also be found for $l<0$. They are symmetric with respect to the case $l\geq 0$ (see the proof of Theorem 2 in Appendix A). However, to solve the issue of identifiability, hereafter we consider the condition $l\geq 0$ as in Theorem 2.
If $l=\frac{1}{2}$ then the sufficient and necessary conditions of the ${\cal GW}_{\nu,\mu,a}$ model in (6) are obtained as special case. It is apparent that, to solve the identifiability problem, $l$ should be fixed (as in the ${\cal GW}$ model) or should depend on the other parameters and/or the dimension, so an interesting question is: which is the “best” $l\geq 0$ in the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ family?
@@
This question can be responded by establishing an optimality criterion. Here we consider the criterion of maximum integral range that can be attained by a compactly supported correlation for a given smoothness and compact support. For an isotropic correlation $\phi\in\Phi_{d}$, the integral range is defined as
$A=\int_{\mathbb{R}^{d}}\phi(\|\boldsymbol{h}\|)\text{d}\boldsymbol{h}=\frac{\pi^{\frac{d}{2}}}{2^{d-1}\Gamma(\frac{d}{2})}\int_{\mathbb{R}^{+}}x^{d-1}\phi(x)\text{d}x.$ (16)
$A$ is the isotropic version of the integral range proposed in Lantuéjoul, (1991) for stationary correlation functions. This measure has a close relation with the ergodicity of the RF, with the definition of an effective number of independent observations, and with the definition of long-range dependence (Yaglom,, 1987; Matheron,, 1989). In particular, following Lim and Teo, (2009), a RF with isotropic correlation $\phi$ is long-range dependent if $A=\infty$ and short-range dependent when $A<\infty$. If $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$ (in particular, if it is compactly supported), then $A=\widehat{\phi}(0)$. Hristopoulos, (2020) interprets $A^{1/d}$ as an average measure of the distance over which two RF values are correlated.
When $d=1$, the integral range corresponds to the area under the unidimensional correlation function. A parametric correlation model that can cover a wide range of areas, from 00 to a specific upper bound, can be considered a flexible model (see Figure  1 and Theorem 6 hereunder). As a consequence, this can serve as a criterion for selecting between two different parametric correlation models, particularly when comparing two compactly supported correlation models that share the same smoothness and compact support parameters. In such cases, the correlation model that achieves a larger integral range is preferred.
@@
For a parametric correlation model $\phi_{\boldsymbol{\theta}}\in\Phi_{d}$, we will denote its associated integral range as $A_{\boldsymbol{\theta},d}$. For instance, using formula 6.561.16 of Gradshteyn and Ryzhik, (2007), for the ${\cal MT}_{\nu,\alpha}$ correlation one has:
@@
The following theorem provides an expression for the integral range associated with the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ correlation model.
@@
For an arbitrary $l\geq 0$, the maximum of $A_{\kappa,\mu,l,a,d}$ is attained when
@@
    $\mu=\mu_{1}(l):=\delta-l+\frac{1}{2}\quad$ if $\quad 0\leq l\leq\frac{d}{2}+\kappa$
@@
    $\mu=\mu_{2}(l):=\sqrt{2\kappa+l^{2}+d+1}-l\quad$ if $\quad l>\frac{d}{2}+\kappa$.
For instance, let $\kappa\geq 0$. If $l=\frac{1}{2}$, then from (13) $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+1}{2},a}={\cal GW}_{\kappa,\mu,a}$ and the maximum integral range is attained when $\mu=\mu_{1}(\frac{1}{2})=\delta=\frac{d+1}{2}+\kappa$, which is the lower bound of ${\cal GW}$ model (6).
@@
The maximum of $A_{\kappa,\mu_{i}(l),l,a,d}$ for $i=1,2$ is attained when $l=d/2+\kappa$.
@@
Accounting for (15), the model with maximum integral range (obtained when $l=\frac{d}{2}+\kappa$ and $\mu=1$), is $\mathcal{GH}_{\delta,\delta+\frac{1}{2},2\delta,a}$ and coincides with the so-called Euclid’s hat correlation model (Gneiting,, 1999) when $\kappa=0$ (triangular or tent model if $d=1$, circular model if $d=2$, spherical model if $d=3$, pentaspherical model if $d=5$, see Section 5) and with an upgraded Euclid’s hat when $\kappa\neq 0$ (Emery and Alegría,, 2022), obtained from the Euclid’s hat by a (fractional) montée or upgrading, as defined by Matheron, (1965, I.4.18).
Hereafter, we call the correlation model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ with $l=\frac{d}{2}+\kappa$ as hypergeometric correlation and we denote as
$\mathcal{H}_{\kappa,\mu,a,d}(x):=\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+d}{2}+\kappa,a}(x),\quad x\geq 0,$ (20)
which is defined as
$\begin{split}\mathcal{H}_{\kappa,\mu,a,d}(x)&=\frac{\Gamma(\kappa+\frac{\mu+1}{2})\Gamma(2\kappa+\frac{d+\mu+1}{2})}{\Gamma(\mu+\frac{d+1}{2}+2\kappa)\Gamma(\kappa+\frac{1}{2})}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\mu+\frac{d-1}{2}+2\kappa}\\ &\times{}_{2}F_{1}\left(\frac{\mu}{2};\frac{\mu+d}{2}+\kappa;\mu+\frac{d+1}{2}+2\kappa;\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\right),\quad x\geq 0.\end{split}$ (21)
Necessary and sufficient validity conditions for $\mathcal{H}_{\kappa,\mu,a,d}$ can be easily obtained from Theorem 2 by setting $l=\frac{d}{2}+\kappa$. Specifically, given the smoothness parameter $\kappa>-\frac{1}{2}$, $\mathcal{H}_{\kappa,\mu,a,d}\in\Phi_{d}$ if and only if $\mu\geq 1$. Interestingly, the conditions are much simpler than the conditions (6) of the $\mathcal{GW}_{\kappa,\mu,a}$ model. In particular, the condition on $\mu$ does not depend on the dimension. The price to pay is that, unlike the $\mathcal{GW}_{\kappa,\mu,a}$ model, the expression of $\mathcal{H}_{\kappa,\mu,a,d}$ depends on the dimension $d$.
The following theorem proves that the ${\cal H}_{\kappa,1,a,d}$ subclass of hypergeometric correlation functions with maximum integral range allows recovering any member of the ${\cal H}_{\kappa,\mu,a,d}$ class, as well as great part of the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ class, by scale mixtures. The chosen optimality criterion (maximizing the integral range) therefore coincides with the greatest flexibility from a modeling viewpoint.
@@
Let $a>0$, $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $\mu>1$ and $l\geq 0$ such that $\mu+2l>1+d+2\kappa$, so that $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}$ owing to Theorem 2. Then, $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ is a beta-type scale mixture of hypergeometric correlation functions of the form $\mathcal{H}_{\kappa,1,b,d}$ with $0<b\leq a$:
$\begin{split}&\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}(x)=\frac{\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu-d}{2}+l)}{\Gamma(\delta+\frac{1-d}{2})\Gamma(\delta+\frac{1}{2}+\kappa)\Gamma(\frac{\mu-1}{2})\Gamma(\frac{\mu-1-d}{2}-\kappa+l)}\\ &\times\int_{0}^{1}\int_{0}^{1}u^{\delta-\frac{d+1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{\delta-\frac{1}{2}+\kappa}(1-v)^{\frac{\mu-d-3}{2}-\kappa+l}\mathcal{H}_{\kappa,1,a\sqrt{uv},d}(x){\rm d}u{\rm d}v,\quad x\geq 0.\end{split}$ (22)
In particular, for $l=\frac{d}{2}+\kappa$, one obtains
$\begin{split}\mathcal{H}_{\kappa,\mu,a,d}(x)&=\frac{\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu}{2}+\kappa)}{\Gamma(\delta+\frac{1-d}{2})\Gamma(\delta+\frac{1}{2}+\kappa)\Gamma^{2}(\frac{\mu-1}{2})}\\ &\times\int_{0}^{1}\int_{0}^{1}u^{\delta-\frac{d+1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{\delta-\frac{1}{2}+\kappa}(1-v)^{\frac{\mu-3}{2}}\mathcal{H}_{\kappa,1,a\sqrt{uv},d}(x){\rm d}u{\rm d}v,\quad x\geq 0.\end{split}$

### v2

@@ 4 Fixing the identifiability problem: the Hypergeometric correlation model
In this section we first refine the validity conditions of the $\mathcal{GH}$ model, clarifying the relationship between necessary and sufficient constraints. We then address the identifiability issue discussed in the previous section by introducing a parsimonious subclass selected according to an explicit optimality criterion.
To this end, we adopt a $\cal GW$}-type reparameterization of the $\mathcal{GH}_{\delta,\beta,\gamma,a}$ model and set
$\delta=\frac{d+1}{2}+\kappa,\qquad\beta=\delta+\frac{\mu}{2},\qquad\gamma=\beta+l.$ (11)
As noted in (9), this parametrization includes the $\cal GW$} model as the special case $l=\frac{1}{2}$.
The following theorem improves Theorem 1 in Emery and Alegría (2022) by providing sharp validity conditions for $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$.
@@
Let $d$ be fixed, $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $a>0$, and $l\geq 0$. Then:
@@
    If $0\leq l\leq\frac{d}{2}+\kappa$, then
    $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}\quad\text{if and only if}\quad\mu\geq\delta-l+\frac{1}{2}.$
@@
    If $l>\frac{d}{2}+\kappa$, then
    $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}\quad\text{if}\quad\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l,$
    and only if
    $\mu\geq\delta-l+\frac{1}{2}.$
For $l=\frac{1}{2}$, the necessary and sufficient validity conditions for $\mathcal{GW}_{\kappa,\mu,a}\in\Phi_{d}$ discussed in Section 2 are recovered as a special case. Specifically, since $l=\frac{1}{2}\leq\frac{d}{2}+\kappa$ if and only if $\kappa\geq\frac{1}{2}-\frac{d}{2}$, which holds for all $d\geq 2$ and $\kappa>-\frac{1}{2}$, or for $d=1$ and $\kappa\geq 0$, Regime 1 of Theorem 2 gives the necessary and sufficient condition
$\mu\geq\frac{d+1}{2}+\kappa,$
For $d=1$ and $\frac{1}{2}<\kappa<0$, instead $l=\frac{1}{2}>\frac{d}{2}+\kappa$, so Regime 2 applies, yielding the sufficient condition $\mu\geq 0.5(\sqrt{8\kappa+9}-1)$ and the necessary condition $\mu\geq\kappa+1$, consistently with the discussion in Section 2.
As discussed in Section 3, the model $\mathcal{GH}_{\delta,\beta,\gamma,a}$ is invariant under the interchange $\beta\leftrightarrow\gamma$, corresponding to $l\mapsto-l$ in the model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$. The admissible region is therefore symmetric with respect to $l=0$, and restricting attention to $l\geq 0$ entails no loss of generality.
Translating Theorem 2 back to the original parameterization $\mathcal{GH}_{\delta,\beta,\gamma,a}$, the following equivalent conditions hold for $\delta>d/2$ and $a>0$:
  1. 1.
    If $0\leq\gamma-\beta\leq\delta-\frac{1}{2}$, then
    $\mathcal{GH}_{\delta,\beta,\gamma,a}\in\Phi_{d}\quad\text{if and only if}\quad\beta+\gamma\geq 3\delta+\frac{1}{2}.$
  2. 2.
    If $\gamma-\beta>\delta-\frac{1}{2}$, then
    $\mathcal{GH}_{\delta,\beta,\gamma,a}\in\Phi_{d}\quad\text{if}\quad\beta+\gamma\geq 2\delta+\sqrt{(\gamma-\beta)^{2}+2\delta},$
    and only if
    $\beta+\gamma\geq 3\delta+\frac{1}{2}.$
In particular, in Regime 1 ($0\leq\gamma-\beta\leq\delta-\frac{1}{2}$), condition (c) of Section 3 implies condition (b), so that (a) and (c) alone characterize validity. Hence, within this regime, the sufficient conditions of Emery and Alegría (2022) reduce to (a) and (c) and are therefore also necessary. In Regime 2 ($\gamma-\beta>\delta-\frac{1}{2}$), whether the gap between sufficiency and necessity can be closed remains an open question.
We now turn to the structural identifiability issue for the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ model. To resolve it, the parameter $l$ must either be fixed (as in the $\cal GW$} model) or linked to the remaining parameters and/or to the spatial dimension. This naturally raises the question: which choice of $l\geq 0$ is optimal ?
@@
We address this question by introducing an explicit optimality criterion that selects a single parsimonious representative once smoothness and compact support are fixed. We adopt the criterion of maximum integral range attainable by a compactly supported correlation with given smoothness and support.
For an isotropic correlation $\phi\in\Phi_{d}$, the integral range is
$A=\int_{\mathbb{R}^{d}}\phi(\|\boldsymbol{h}\|)\,\mathrm{d}\boldsymbol{h}=\frac{\pi^{\frac{d}{2}}}{2^{d-1}\Gamma(\frac{d}{2})}\int_{0}^{\infty}x^{d-1}\phi(x)\,\mathrm{d}x.$ (12)
This quantity, originally introduced by Lantuéjoul (1991), provides a global and threshold-free measure of dependence (Yaglom, 1987; Matheron, 1989; Hristopulos and Zukovic, 2011). If $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$ (in particular, if it is compactly supported), then $A=\widehat{\phi}(0)$, so that the integral range coincides with the spectral density at zero frequency. When $d=1$, AA reduces to the area under the correlation function.
For compactly supported correlations, maximizing $A$ therefore selects, among all admissible models with fixed smoothness and support, the one with maximal low-frequency spectral power, that is, the strongest large-scale dependence compatible with the imposed constraints. While alternative optimality principles could be considered, they would typically depend on the sampling design or on a specific loss function. In contrast, the integral range criterion is intrinsic to the model class and yields a reproducible, one-to-one specification within the family.
@@
For a parametric model $\phi_{\boldsymbol{\theta}}\in\Phi_{d}$, we denote the associated integral range by $A_{\boldsymbol{\theta},d}$. For example, for the Matérn correlation one has
@@
The following theorem provides the integral range of $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$.
@@
Let $\kappa>0$, $a>0$, and $d\in\mathbb{N}$ be fixed. For each $l\geq 0$, the maximum of $A_{\kappa,\mu,l,a,d}$ with respect to $\mu$ (over its admissible parameter space) is attained at $\mu^{\star}(l)$, where:
@@
    if $0\leq l\leq\dfrac{d}{2}+\kappa$, then
    $\mu^{\star}(l)=\mu_{1}(l):=\delta-l+\frac{1}{2};$
@@
    if $l>\dfrac{d}{2}+\kappa$, then
    $\mu^{\star}(l)=\mu_{2}(l):=\sqrt{2\kappa+l^{2}+d+1}-l.$
For instance, let $\kappa\geq 0$. If $l=\frac{1}{2}$, then from (9) $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+1}{2},a}={\cal GW}_{\kappa,\mu,a}$ and the maximum integral range is attained when $\mu=\mu_{1}(\frac{1}{2})=\delta=\frac{d+1}{2}+\kappa$, which is the lower bound of $\cal GW$} model in this case.
@@
Let $\kappa>-1/2$, $a>0$, and $d\in\mathbb{N}$ be fixed, and set $\delta=\frac{d}{2}+\kappa$.
  1. 1.
    The function $l\mapsto A_{\kappa,\mu_{1}(l),l,a,d}$ is non-decreasing on $0,\delta]$, hence
    $\max_{0\leq l\leq\delta}A_{\kappa,\mu_{1}(l),l,a,d}=A_{\kappa,\mu_{1}(\delta),\delta,a,d}=A_{\kappa,1,\delta,a,d}.$
  2. 2.
    The function $l\mapsto A_{\kappa,\mu_{2}(l),l,a,d}$ is non-increasing on $(\delta,\infty)$ and
    $\lim_{l\downarrow\delta}A_{\kappa,\mu_{2}(l),l,a,d}=A_{\kappa,1,\delta,a,d}.$
In particular, the global maximum over all admissible $(\mu,l)$ is attained at $(\mu,l)=(1,\delta)$ and is equal to $A_{\kappa,1,\delta,a,d}$.
@@
Accounting for (11), the model with maximum integral range (obtained when $l=\frac{d}{2}+\kappa$ and $\mu=1$), is $\mathcal{GH}_{\delta,\delta+\frac{1}{2},2\delta,a}$ and coincides with the so-called Euclid’s hat correlation model (Gneiting, 1999) when $\kappa=0$ (triangular or tent model if $d=1$, circular model if $d=2$, spherical model if $d=3$, pentaspherical model if $d=5$, see Section 5) and with an upgraded Euclid’s hat when $\kappa\neq 0$ (Emery and Alegría, 2022), obtained from the Euclid’s hat by a (fractional) montée or upgrading, as defined by Matheron (1965, I.4.18).
Fixing $l=\frac{d}{2}+\kappa$, we define the hypergeometric correlation by
$\mathcal{H}_{\kappa,\mu,a,d}(x):=\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+d}{2}+\kappa,a}(x),\qquad x\geq 0.$ (16)
An explicit expression is given by
$\begin{split}\mathcal{H}_{\kappa,\mu,a,d}(x)&=\frac{\Gamma\!\left(\kappa+\frac{\mu+1}{2}\right)\Gamma\!\left(2\kappa+\frac{d+\mu+1}{2}\right)}{\Gamma\!\left(\mu+\frac{d+1}{2}+2\kappa\right)\Gamma\!\left(\kappa+\frac{1}{2}\right)}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\mu+\frac{d-1}{2}+2\kappa}\\[4.0pt] &\quad\times{}_{2}F_{1}\!\left(\frac{\mu}{2},\frac{\mu+d}{2}+\kappa;\mu+\frac{d+1}{2}+2\kappa;\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\right),\qquad x\geq 0.\end{split}$ (17)
Notably, the model $\mathcal{H}_{\kappa,\mu,a,d}$ always lies in Regime 1 of Theorem 2, where validity is completely characterized. Indeed, by setting $l=\frac{d}{2}+\kappa$ in Theorem 2, we obtain the condition
$\mathcal{H}_{\kappa,\mu,a,d}\in\Phi_{d}\quad\text{if and only if}\quad\mu\geq 1.$
These validity conditions are substantially simpler than those for the $\mathcal{GW}_{\kappa,\mu,a}$ model discussed in Section 2. In particular, the lower bound on $\mu$ is independent of the spatial dimension. The trade-off is that, unlike the $\mathcal{GW}_{\kappa,\mu,a}$ model, the expression of $\mathcal{H}_{\kappa,\mu,a,d}$ explicitly depends on $d$.
The next theorem shows that the subclass ${\cal H}_{\kappa,1,a,d}$, obtained by maximizing the integral range, retains substantial modeling flexibility. Through beta-type scale mixtures, it can generate any member of ${\cal H}_{\kappa,\mu,a,d}$ and a broad subclass of the Gauss–hypergeometric family $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$. Thus, the maximal integral range criterion does not restrict modeling capacity, but instead selects a parsimonious representative that remains rich within the admissible class.
@@
Let $a>0$, $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $\mu>1$, and $l\geq 0$ such that $\mu+2l>1+d+2\kappa$, so that $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}$ by Theorem 2. Then $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ admits a beta-type scale mixture representation in terms of hypergeometric correlations $\mathcal{H}_{\kappa,1,b,d}$ with $0<b\leq a$:
$\begin{split}\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}(x)&=\frac{\Gamma\!\left(\delta+\frac{\mu-d}{2}\right)\Gamma\!\left(\delta+\frac{\mu-d}{2}+l\right)}{\Gamma\!\left(\delta+\frac{1-d}{2}\right)\Gamma\!\left(\delta+\frac{1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu-1}{2}\right)\Gamma\!\left(\frac{\mu-1-d}{2}-\kappa+l\right)}\\ &\quad\times\int_{0}^{1}\!\!\int_{0}^{1}u^{\delta-\frac{d+1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{\delta-\frac{1}{2}+\kappa}(1-v)^{\frac{\mu-d-3}{2}-\kappa+l}\\ &\qquad\qquad\times\mathcal{H}_{\kappa,1,a\sqrt{uv},d}(x)\,\mathrm{d}u\,\mathrm{d}v,\qquad x\geq 0.\end{split}$ (18)
In particular, when $l=\frac{d}{2}+\kappa$, one obtains
$\begin{split}\mathcal{H}_{\kappa,\mu,a,d}(x)&=\frac{\Gamma\!\left(\delta+\frac{\mu-d}{2}\right)\Gamma\!\left(\delta+\frac{\mu}{2}+\kappa\right)}{\Gamma\!\left(\delta+\frac{1-d}{2}\right)\Gamma\!\left(\delta+\frac{1}{2}+\kappa\right)\Gamma^{2}\!\left(\frac{\mu-1}{2}\right)}\\ &\quad\times\int_{0}^{1}\!\!\int_{0}^{1}u^{\delta-\frac{d+1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{\delta-\frac{1}{2}+\kappa}(1-v)^{\frac{\mu-3}{2}}\\ &\qquad\qquad\times\mathcal{H}_{\kappa,1,a\sqrt{uv},d}(x)\,\mathrm{d}u\,\mathrm{d}v,\qquad x\geq 0.\end{split}$